On October 1, 2013, Ezzy Corporation issued a 6-year bond worth $387,000 with an interest rate of 10% per annum. Interest is to be paid semi-annually on March 31 and September 30. At the time of the issuance, the market interest rate was 6%. Ezzy Corporation amortizes any premium or discount using the straight-line method.

Calculate the bond issue price and the resulting premium or discount.

Do not enter dollar signs or commas in the input boxes.
Round your answers to 2 decimal places.
Use the present value tables in the textbook.

Bond issue price: $Answer

Premium or discount: $Answer
Answer #1

Bond issue price = Present value of interest+Present value of maturity

                          = (387000*5%*9.954)+(387000*0.70138)

Bond issue price = $464043.96

Premium = 464043.96-387000 = 77043.96